An aluminum and magnesium allow weighs 26kg. The ratio of aluminum to magnesium is 8:5. How many kilograms of magnesium are in the alloy
Answer by ikleyn(52830)

.

We have   8 + 5 = 13  equal parts  (by mass).

Each part is   26/13 = 2 kilograms.

Hence, there are   8*2 = 16 kg aluminum and   5*2 = 10 kg magnesium.          ANSWER

Solved / mentally.